OpenStudy (anonymous):

Let tan x=12 and cot y=16. What is cot X + tan Y ?

OpenStudy (raden):

use these identity : tan = 1/cot cot = 1/tan

OpenStudy (raden):

if tan x = 12 then cot x = 1/tan x = 1/12 while if cot y = 16 then tan y = 1/cot y = 1/16

OpenStudy (raden):

so, cot x + tan y = 1/12 + 1/16 = ... ?

OpenStudy (anonymous):

0.1458

OpenStudy (anonymous):

is that correct

OpenStudy (raden):

1/12 + 1/16 = 4/48 + 3/48 = 7/48 = 0.1458 you are right

OpenStudy (anonymous):

Thank you so, so much very grateful !!!!! ;D

OpenStudy (raden):

welcome :)